Song earns $282,500 taxable income as an interior designer and is taxed at an average rate of 20 percent (i.e., $56,500 of tax).

   

a. If Congress increases the income tax rate such that Song’s average tax rate increases from 20 percent to 25 percent, how much more income tax will she pay assuming that the income effect is descriptive? (Round your intermediate calculations and final answer to 2 decimal places.)

    

b. If the income effect is descriptive, the tax base and the tax collected will increase.

True
False
Answer #1

a)

After-tax income =282500-56500 = 226,000

Taxable income at 25% rate:

=226000/75% = 301,333,333

Tax = 301,333,333-226000= 75333.3333

Extra tax = 75333.333-56500= 18833.33

b)

True.